OpenStudy (walters):

Let G be a cyclic group with generator a, |a|=20 and H=.Find the order of a^3H in G/H.

OpenStudy (walters):

\[G=\left\{ e,a ^{1},a ^{2},a ^{3},a ^{4},a ^{5},... a ^{19}\right\}\]

OpenStudy (walters):

i found that\[<a ^{4}>=\left\{ e,a ^{4},a^{8},a ^{12},a ^{16} \right\}\] This means \[a ^{3}<a ^{4}>=\left\{ a ^{3},a ^{7},a ^{11},a ^{15},a ^{19} \right\}=a ^{3}H\]

OpenStudy (walters):

THEN WHAT WILL BE THE ORDER OF \[a ^{3}H\]

OpenStudy (walters):

@terenzreignz

terenzreignz (terenzreignz):

is any element of a^3 (except e) in the subgroup <a^4> ?

OpenStudy (walters):

no

terenzreignz (terenzreignz):

of course there is :) a^12

OpenStudy (walters):

not all of them are in <a^4>

terenzreignz (terenzreignz):

just one is needed... a^3 <a^4> a^6 <a^4> a^9 <a^4> a^12 <a^4> = <a^4> order is 4.

OpenStudy (walters):

i thought they want the smallest integer such that a^n =e

terenzreignz (terenzreignz):

that would be the order of a, but that's not what they're asking :) They're asking the smallest integer such that a^3n <a^4> = <a^4>

OpenStudy (walters):

so wat would be the difference if they have said find the order ofa^3H

terenzreignz (terenzreignz):

big difference. as was demonstrated, the order of a^3 <a^4> is 4

OpenStudy (walters):

G/H={H,aH,a^2H,a^3H} so wat is the purpose of ths statement on this question

OpenStudy (walters):

so it means even if they have said . Find the order of a^3H we will get the same answer

OpenStudy (walters):

by excluding in G/H

terenzreignz (terenzreignz):

we're just finding the the order of a^3H considering that <a^3 H> is a subgroup of G/H

OpenStudy (walters):

OK :)

OpenStudy (walters):

DID U DO THE EQUIVALENCE RELATION